OpenStudy (anonymous):

Solve: –4(2x – 1) = –6(x + 2) – 2

OpenStudy (anonymous):

–4(2x – 1) = –6(x + 2) – 2 -8x+4=-6x-12-2 -8x+4=-6x-14 -2x=-18 x=9

OpenStudy (anonymous):

okay..? u would distribute then try to isolate x